### v1

@@ 7 Experiments and Results
    The efficient structures of feature graphs (Section 7.2)
@@
    The dependency between edges and interaction pairs (Section 7.3)
@@
    The relationship between multi-hops and number of message passing layers (Section 7.4)
@@
    Scalability of efficient feature graphs (Section 7.5)
Since Section 7.2 - 7.4 consider the variants of graph structures w.r.t. the given dataset, we show only results from an equation containing 2 interaction terms and 2 unary terms, i.e. $p=2,q=2$. In Section 7.5, $p$ is varied while $q=2$.
7.1 Experimental Setting
@@
Each synthetic dataset contains 10,000 samples generated by distribution described in Methodology section (Section 6). In every randomization, we use seed number $n$ for feature $x_{n}$ and seed number 0 for noise term. We split it into first 7,000 rows for training and next 3,000 rows for testing sets.
@@
7.2 Interaction Edges vs. Non-interaction Edges
@@
7.3 Removing Interaction Edges from a Graph
@@
The lower-sided confidence interval 95% is shown in Table 2. It ensures us that interaction edges are significantly important for input feature graph in attention-based GNNs. Figure 6 shows the MAE of some random sample generated, sorted by the MAE of $E(G)\setminus\{\{0,1\},\{2,3\}\}$ from the 3 message passing layer model to illustrate how they are.
@@
Table 2: 95% lower-sided confidence interval of paired-difference of MAE between a graph and interaction-edge-removed graph
@@
7.4 Reachability of Interaction Nodes and the Number of Message Passing Layers through a Longer Path
@@
7.5 Limitation of Complete Feature Graphs
@@ 7.6 Discussion
However, the results shown in this section are from only one simulated dataset. Even though the results seem obviously to infer the conclusions, we need some theoretical results to guarantee the consistency of the results in any dataset that we are focusing on. The results from Sections 7.2 and 7.3 motivate us to reformulate the problem into a theoretical perspective which is discussed in the next section.
8 Case Study: Click-through-rate Prediction
@@
In this experiment, we utilized the Criteo dataset [25] to evaluate the performance of different graph structures using the AUC score as the primary metric.
FiGNN, a widely recognized graph model for CTR prediction, achieves an AUC score of 0.8062. Initially, we generated random sparse graphs with varying percentages of connections and evaluated their performance using a simple GNN model. The optimal AUC score for randomly constructed graphs was approximately 0.8072, achieved with 24% graph connections. Subsequently, we experimented with a fully connected graph, utilizing 100% of the graph connections. The AUC score for the fully connected graph was 0.8073. Additionally, we manually constructed a sparse graph inspired by the GDCN heatmap [26]. This manually constructed graph achieved an AUC score of 0.8076 with only 13 % of the connections.
The findings of this experiment demonstrate that a simple GNN model with a randomly constructed sparse graph can outperform the FiGNN model and achieve results comparable to a fully connected graph. Moreover, the manually constructed sparse graph, designed with deliberate and meaningful connections, outperforms both the fully connected graph and FiGNN in performance. This supports the results found in our simulation.
9 Conclusion and Future Work
In this study, we investigated how the structure of feature graphs relates to their capacity to represent pairwise feature interactions within GNNs. We offered both theoretical and empirical insights on the development of sparse feature graphs. Our findings reveal that interaction edges are crucial, whereas non-interaction edges can introduce noise, which hinders the learning process of GNN models. From a theoretical standpoint, employing the Minimum Description Length (MDL) principle, we showed that feature graphs that retain only essential interaction edges provide a more efficient and interpretable representation compared to complete graphs. Empirical tests on synthetic datasets underscored the significance of interaction edges for improved prediction performance, while non-interaction edges were found to introduce noise that reduces model accuracy. Additionally, we discovered that the connection between interacting nodes, even without direct interaction edges, can enhance learning via multi-hop message passing in GNNs.
Our findings emphasize the significance of feature graph sparsity for both computational efficiency and prediction accuracy, particularly in large-scale datasets where complete graphs become useless. Furthermore, we validated the idea of our insights through a case study on click-through rate prediction, further highlighting the practical value of constructing well-designed feature graphs for real-world tasks.
@@ 9 Conclusion and Future Work
    Design new message-passing mechanisms specifically tailored for interaction modeling, enabling GNNs to better capture interaction patterns and relationships in feature graphs.
By addressing these directions, we hope to further advance the understanding and utility of feature graphs in GNN-based models, contributing to improved model performance and interpretability across a broad range of applications.

### v2

@@ VII Experiments and Results
    The efficient structures of feature graphs (Section VII-B)
@@
    The dependency between edges and interaction pairs (Section VII-C)
@@
    The relationship between multi-hops and number of message passing layers (Section VII-D)
@@
    Scalability of efficient feature graphs (Section VII-E)
Since Section VII-B - VII-D consider the variants of graph structures w.r.t. the given dataset, we show only results from an equation containing 2 interaction terms and 2 unary terms, i.e. $p=2,q=2$. In Section VII-E, $p$ is varied while $q=2$.
VII-A Experimental Setting
@@
Each synthetic dataset contains 10,000 samples generated by distribution described in Methodology section (Section VI). In every randomization, we use seed number $n$ for feature $x_{n}$ and seed number 0 for noise term. We split it into first 7,000 rows for training and next 3,000 rows for testing sets.
@@
VII-B Interaction Edges vs. Non-interaction Edges
@@
VII-C Removing Interaction Edges from a Graph
@@
The lower-sided confidence interval 95% is shown in Table II. It ensures us that interaction edges are significantly important for input feature graph in attention-based GNNs. Figure 6 shows the MAE of some random sample generated, sorted by the MAE of $E(G)\setminus\{\{0,1\},\{2,3\}\}$ from the 3 message passing layer model to illustrate how they are.
@@
TABLE II: 95% lower-sided confidence interval of paired-difference of MAE between a graph and interaction-edge-removed graph
@@
VII-D Reachability of Interaction Nodes and the Number of Message Passing Layers through a Longer Path
@@
VII-E Limitation of Complete Feature Graphs
@@ VII-F Discussion
However, the results shown in this section are from only one simulated dataset. Even though the results seem obviously to infer the conclusions, we need some theoretical results to guarantee the consistency of the results in any dataset that we are focusing on. The results from Sections VII-B and VII-C motivate us to reformulate the problem into a theoretical perspective which is discussed in the next section.
VIII Case Study: Click-through-rate Prediction
@@
In this experiment, we utilized the Criteo dataset [31] to evaluate the performance of different graph structures using the AUC score as the primary metric. We adopt FiGNN, a widely recognized graph model for CTR prediction, as our baseline. FiGNN utilizes the complete graph and learns the dynamic graph structure for each sample, achieving an AUC score of 0.8062.
Initially, we generated random sparse graphs with varying percentages of connections and evaluated their performance using a simple GNN model. The optimal AUC score for randomly constructed graphs was approximately 0.8072, achieved with 24% graph connections. Subsequently, we experimented with a fully connected graph, utilizing 100% of the graph connections. The AUC score for the fully connected graph was 0.8073. Additionally, we manually constructed a sparse graph inspired by the GDCN heatmap [32]. This manually constructed graph achieved an AUC score of 0.8076 with only 13 % of the connections.
The findings of this experiment demonstrate that a simple GNN model with a randomly constructed sparse graph can outperform the FiGNN model and achieve results comparable to a fully connected graph. Moreover, the manually constructed sparse graph, designed with connections that align with the feature interactions, outperforms both the fully connected graph and FiGNN in performance.
IX Conclusion and Future Work
In this study, we investigated the relationship between the structure of feature graphs and their capacity to represent pairwise feature interactions within GNNs. We offered both theoretical and empirical insights on the development of sparse feature graphs. Our findings reveal that interaction edges are crucial, whereas non-interaction edges can introduce noise, which hinders the learning process of GNN models. From a theoretical standpoint, employing the Minimum Description Length (MDL) principle, we showed that feature graphs that retain only essential interaction edges provide a more efficient and interpretable representation compared to complete graphs. Empirical tests on synthetic datasets underscored the significance of interaction edges for improved prediction performance, while non-interaction edges were found to introduce noise that reduces model accuracy. Additionally, we discovered that the connection between interacting nodes, even without direct interaction edges, can enhance learning via multi-hop message passing in GNNs.
Our findings emphasize the significance of feature graph sparsity for both computational efficiency and prediction accuracy, particularly in large-scale datasets where complete graphs become useless. Furthermore, we validated our insights through a case study on click-through rate prediction, further highlighting the practical value of constructing well-designed feature graphs for real-world tasks.
@@ IX Conclusion and Future Work
    Design new message-passing mechanisms specifically tailored for interaction modeling, enabling GNNs better to capture interaction patterns and relationships in feature graphs.
By addressing these directions, we aim to enhance the understanding and utility of feature graphs in GNN-based models, thereby contributing to improved model performance and explainability across a wide range of applications.
Acknowledgment
Anonymous submission.